# GastCoCo: Graph Storage and Coroutine-Based Prefetch Co-Design for Dynamic Graph Processing

### v3

@@ 7.6. Real-World Graph Updates
Firstly, we conduct a properties update experiment using the OGB dataset. The results are shown in Figure 16. Since updating properties does not cause structural changes, many systems have shown good performance in this task. For graph data with large properties, using AOE in CBList is more beneficial for updates because each edge’s property has better spatial locality with the destination vertex ID. Conversely, using AOA is more beneficial for graph algorithms that do not require properties (e.g., PageRank), as the spatial locality between the destination vertex IDs of neighboring edges is better. The conclusion can also be confirmed by our experiment shown in Figure 18, where AOA has a lower runtime than AOE in graph algorithms (i.e., PageRank), while AOE achieves higher throughput in update performance.
@@
7.7. Trade-off between Throughput and Latency
GastCoCo supports both batch and individual edge updates. The granularity of batch updates can be adjusted according to specific applications. Larger batch sizes can lead to higher throughput, which means that each edge must wait for more edges to be updated, resulting in data updates with lower freshness. To explain this, we conduct an experiment to measure the trade-off between throughput and data freshness, as shown in Figure 18.
As the batch size increases, throughput rises, but latency also increases. Additionally, at larger batch sizes, the increase in throughput becomes marginal. From Figure 18, it is evident that when the batch size is $10^{4}$ or $10^{5}$, GastCoCo achieves high throughput while maintaining low latency. Therefore, unless the upper-layer application specifies a fixed batch size, GastCoCo will choose a batch size of $10^{4}$ or $10^{5}$ when actively initiating batch update mode.
@@ 7.8. Performance Analysis
We conduct performance analysis experiments on SSSP and batch insertions for our designs to understand their performance gains and losses. SSSP includes GetVertices($cond$), thus increasing the probability of hardware prefetching failures, making software prefetching more effective in comparison. Batch insertions utilize update/query pointers in CBList for locating update positions within the edge storage structure, hence hybrid strategy I cannot be applied as it is only implementable on GTChain. In particular, Sequential Execution (SE) means executing tasks in sequential execution mode (introduced in subsection 5.1); Interleaved Execution (IE) means using coroutines to form an interleaved execution mode (introduced in subsection 5.1) during computation; Software Prefetching (SP) means incorporating software prefetching instructions during the interleaved execution process; Hybrid I and Hybrid II represent employing hybrid strategy I with hardware failure time and hybrid strategy II with the content to be prefetched proposed in subsection 6.2.
From Figure 14, we could observe that the interleaved execution mode (IE no SP) spends a longer running time compared with the sequential execution mode (SE) due to the additional switching cost of coroutines. However, if we further employ software prefetching and hybrid prefetching strategy, GastCoCo achieves at most 1.83×1.83$\times$1.83 × (graph computation) and 1.62×1.62$\times$1.62 × (graph update) speedup compared with sequential execution on CBList.
@@
7.9. Space Cost
Figure 19. Memory consumption (GB) of various systems for storing different graphs.
We conduct a memory consumption experiment, as shown in Figure 19. Although GastCoCo incurs additional overhead due to GTChain and B+ tree indexing, it limits the edge block size to four times the cache line size, reserving less space compared to most other systems. Therefore, in terms of memory consumption, GastCoCo outperforms other systems, except for PCSR and, in some cases, Terrace and Sortledton.
Figure 20. Sensitivity to cache line size multiple, coroutine, and hardware configurations.
7.10. Varying Hardware Environments
We test the impact of system parameters (including cache line size multiple and the number of coroutines) on the performance of GastCoCo on Hollywood across four different hardware environments. In addition to the CPU used in other experiments, we also use three other CPUs: an 8-core Intel Xeon Platinum 8369B, a 16-core Intel Xeon Platinum 8163, and a 32-core Intel Xeon Platinum 8369HC. Since the cache line size multiple cannot be changed after GastCoCo loads the data, its value must balance the performance of both graph computation and graph update tasks. As shown in Figure 20(a) and (b), as the cache line size multiple increases, the chunk size of edge blocks in CBList also increases, resulting in an initial performance improvement followed by a decline. This is because larger chunks initially allow for more compact data storage, enhancing performance. However, excessively large chunk sizes lead to more sparse storage overall, reducing spatial locality and subsequently decreasing performance. Additionally, excessively large chunks can cause prefetching to be incomplete before switching back, leading to failed switches. GastCoCo achieves superior performance in both PageRank and batch update tasks in four hardware environments when the cache line size multiple is set to 4. As shown in Figure 20(c) and (d), as the number of coroutines increases, the performance of PageRank decreases because each computation takes a relatively long time. The more coroutines in the coroutine pool, the longer other coroutines will have to wait, causing the prefetched data to become stale or be replaced. Conversely, in batch update tasks, where computation times are shorter, the initial increase in the number of coroutines does not lead to cache data becoming stale. However, as the number of coroutines continues to increase, the issue of cache data staleness reappears. The performance of PageRank algorithm is good with 2 coroutines. For batch update tasks, using 16 coroutines achieves better acceleration in most environments. So the number of coroutines per thread is typically set to 2 for graph analysis tasks and around 16 for graph update tasks.
@@ 8. Related Work
Dynamic graph storage systems. There have been graph storage systems developed for dynamic graph storage and processing, e.g., (Pandey et al., 2021; Fuchs et al., 2022; Macko et al., 2015; Dhulipala et al., 2019; Leo and Boncz, 2021; Zhu et al., 2020; Feng et al., 2020; Kumar and Huang, 2020; Wheatman and Xu, 2018; Kyrola et al., 2012). LiveGraph (Zhu et al., 2020) adopts a log-structured design, allocating a contiguous edge block to each vertex for fast neighbor scanning and enabling swift updates through append-only insertions. However, this method leads to unordered neighbor storage, impeding efficient execution of common operations like intersections. Moreover, once fixed-size edge blocks are full, data must be copied to larger blocks to maintain continuity, reducing insertion speed. LLAMA (Macko et al., 2015) uses a multi-version array storage to enhance spatial locality in graph analysis. However, continuous data insertion leads to numerous array versions and scatters a vertex’s neighbors across them, still causing extensive CPU cache stalls during analysis. Aspen (Dhulipala et al., 2019) use a tree structure for edge data to enable fast insertion and maintain ordered storage of neighbors. However, this structure results in minimal inherent spatial locality among data nodes, leading to poor access performance. Terrace (Pandey et al., 2021) and Sortledton (Fuchs et al., 2022) adopt different data structures for storing neighbors of vertices of varying degrees to enhance data locality and minimize cache misses during access. However, for vertices with a high degree, they are still stored in discontinuous structures, and neighbors of a single vertex may span multiple structures, ultimately impacting search and graph algorithm performance. Teseo (Leo and Boncz, 2021) and PCSR (Wheatman and Xu, 2018), employing CSR-like structures, use reserved gaps to prevent severe data movement overhead. However, they still face the same challenges as CSR when facing a surge in data updates. Different from existing systems, we design a novel dynamic graph data structure to facilitate rapid graph insertions. It integrates the advantages of hardware prefetching, software prefetching, and coroutines to minimize cache misses, thereby enhancing graph computation performance.

### v4

@@ 7.6. Real-World Graph Updates
Firstly, we conduct a properties update experiment using the OGB dataset. The results are shown in Figure 16. Since updating properties does not cause structural changes, many systems have shown good performance in this task.
@@ 7.7. Performance Analysis
We conduct performance analysis on SSSP and batch insertions for our designs to understand the performance gains and losses. SSSP includes GetVertices($cond$), making the probability of hardware prefetching failures increase and software prefetching more effective. Batch insertions utilize update/query pointers in CBList for locating update positions in the edge storage structure. Hence, the hybrid strategy I cannot be applied as it is only implementable on GTChain. In particular, Sequential Execution (SE) means executing tasks in sequential execution mode (refer to subsection 5.1); Interleaved Execution (IE) means using coroutines to form an interleaved execution mode (refer to subsection 5.1); Software Prefetching (SP) means incorporating software prefetching instructions during the interleaved execution process; Hybrid I and Hybrid II represent employing hybrid strategy I with hardware failure time and hybrid strategy II with the content to be prefetched respectively in section 6. From Figure 14, we could observe that the interleaved execution mode (IE no SP) spends longer running time compared with the sequential execution mode (SE) due to the additional switching cost of coroutines. However, if we further employ software prefetching and hybrid prefetching strategy, GastCoCo achieves at most 1.83×1.83$\times$1.83 × (graph computation) and 1.62×1.62$\times$1.62 × (graph update) speedup compared with sequential execution on CBList.
@@ 8. Related Work
Dynamic graph storage systems. There have been graph storage systems developed for dynamic graph storage and processing  (Pandey et al., 2021; Fuchs et al., 2022; Macko et al., 2015; Dhulipala et al., 2019; Leo and Boncz, 2021; Zhu et al., 2020; Feng et al., 2020; Kumar and Huang, 2020; Wheatman and Xu, 2018; Kyrola et al., 2012). LiveGraph (Zhu et al., 2020) adopts a log-structured design, allocating a contiguous edge block to each vertex for fast neighbor scanning and enabling swift updates through append-only insertions. However, this method leads to unordered neighbor storage, impeding the efficient execution of common operations, and data must be copied to larger blocks to maintain continuity once current edge blocks are full. LLAMA (Macko et al., 2015) uses multi-version array storage to enhance spatial locality in graph analysis. However, continuous data insertion leads to numerous array versions and scatters a vertex’s neighbors across them, still causing extensive CPU cache stalls during analysis. Aspen (Dhulipala et al., 2019) uses a tree structure for edge data to enable fast insertion and maintain ordered storage of neighbors. However, this structure results in minimal inherent spatial locality among data nodes, leading to poor access performance. Terrace (Pandey et al., 2021) and Sortledton (Fuchs et al., 2022) adopt different data structures for storing neighbors of vertices of varying degrees to enhance data locality and minimize cache misses during access. However, for vertices with a high degree, they are still stored in discontinuous structures, and neighbors of a single vertex may span multiple structures, ultimately impacting search and graph algorithm performance. Teseo (Leo and Boncz, 2021) and PCSR (Wheatman and Xu, 2018), employing CSR-like structures, use reserved gaps to prevent severe data movement overhead. However, they still face the same challenges as CSR when facing a surge in data updates. Different from existing systems, we design a novel dynamic graph data structure to facilitate rapid graph insertions. It integrates the advantages of hardware prefetching, software prefetching, and coroutines to minimize cache misses, thereby enhancing graph computation performance.
@@ 9. Conclusion
Acknowledgements.
This work was supported by the National Key R&D Program of China (2023YFB4503601), the National Natural Science Foundation of China (U2241212, 62072082, and 62202088), the 111 Project (B16009), the Joint Funds of Natural Science Foundation of Liaoning Province (2023-MSBA-078), the Fundamental Research Funds for the Central Universities (N2416011), and a research grant from Alibaba Innovative Research (AIR) Program. Yanfeng Zhang and Shufeng Gong are the co-corresponding authors.